Using any method find the volume of the solid between the surfaces z= 8 – y² and z = 22 + y2
Answer #1

Clue and Given circular parabolaid eportion z = 8-2-2 this Sulia Pruded z = 2 + 4 Dej aduan pemakat? again x² + y 222 +2y = 8where dy is dv is equal to releder.de Im cylindrical.coorralimate volume obedy dz= rardada So ve fff or forte de la And limitaround the circle colo which is from -22010 ardz drdo 27 8-82 Ilean da do 27 8.92 Oz dardo 22 2 ra Tor (8-82) - 83 dodo -23 2